OpenStudy (anonymous):

Model the function rule y = -2x - 3 with a table of values

OpenStudy (anonymous):

all you have to do is choose a value for x and find the y, or vice versa.

OpenStudy (anonymous):

could help cus i still dont get it

OpenStudy (anonymous):

y=-2x-3 y=-2(1)-3 y=-2-3 y=-5 so if x=1, then y=-5

OpenStudy (anonymous):

do this with several numbers, usually I use -2,-1,0,1,2

OpenStudy (anonymous):



OpenStudy (anonymous):

fill out the rest. (:

OpenStudy (anonymous):

loll i cant.....